because of the jet stream, it takes 8 hours to fly from Orlando to London and 9.5 hours to return. if an airplane averages 550 mph in still air, what is the average rate of the wind blowing in the direction from Orlando to London?
Answer by Boreal(15235)

The distance between the two cities doesn't change.
going east, 8(550+w)=distance
going west 9.5(550-w)
set them equal and distribute the constant.
4400+8w=5225-9.5w
17.5w=825
w=47.14 or 47.1 mph average wind speed ANSWER
going over speed is 597.14 mph*8 hours=4777.12 miles
coming back speed is 502.86 mph*9.5 hours=4777.17 miles, rounding error only.